OpenStudy (anonymous):

Plzzz Help! Geometric series: Find a formula for Sn for: √3 + 3 + 3√3 + 9 + ... to n terms.

OpenStudy (anonymous):

\[S _{n}=\frac{ a(1-r ^{n}) }{ 1-r }\]

OpenStudy (anonymous):

a=is the first term in the sequence r=common ratio

OpenStudy (anonymous):

if your common ratio is greater than 1, \[S _{n}=\frac{ a(r ^{n}-1) }{ r-1 }\]

OpenStudy (anonymous):

\[S_{n}=\frac{ \sqrt{3} (\sqrt{3}^{n}-1)}{ \sqrt{3} -1}\] How do I simplify this? The answer is not this.

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

you have the nth term as \(a_n=3^{n/2}\)

OpenStudy (anonymous):

and are you sure the sum is not to infinity?

OpenStudy (anonymous):

the answer is \[S_{n}=\frac{ 3+\sqrt{3} }{ 2 } ((\sqrt{3})^{n} - 1)\]

OpenStudy (anonymous):

they rationalized the denominator. \[\]

OpenStudy (anonymous):

\[\frac{\sqrt{3}}{\sqrt{3}-1}\times{\sqrt{3}+1\over\sqrt{3}+1}=\] simplify

OpenStudy (anonymous):

ok thnx a lot @kausarsalley and @electrokid !